OpenStudy (anonymous):

Calculus III Questions: 1. Let Q = (1, -1, 2), P = (1, 3, -2) and N = <1, 2, 2>. Find the point of the intersection of the line through P in the direction of N, and the plane through Q perpendicular to N. 2. Find the distance between the indicated point and plane. (a) (1, 1, 2) and 3x + y - 5z = 2 (b) (3, -2, 1) and the yz-plane

OpenStudy (amistre64):

you got that one figured out :)

OpenStudy (anonymous):

haha hello again. I was hoping maybe someone could do #1 again. if not, think you could help with #2?

OpenStudy (amistre64):

1 is hopeless til you get the notation cleared up 2 is better

OpenStudy (anonymous):

2 is in the cross product chapter, so I assume cross product use is required

OpenStudy (amistre64):

the distance from a plane is defined perpendicular to it; so we need to use the normal vector from the plane, create a line equation from the point and see where it intersects the plane at; then measure the distance between the points

OpenStudy (amistre64):

write me up the line equations

OpenStudy (anonymous):

the normal for 2a would be <3, 1, -5>

OpenStudy (amistre64):

good

OpenStudy (anonymous):

parametric form: x = 1 + 3t y = 1 + t z = 2 -5t

OpenStudy (amistre64):

now use those values for xyz into the plane equation to find the point it intersects at

OpenStudy (amistre64):

or rather the "t" value for the line

OpenStudy (anonymous):

substitutiing the parametric forms into the equation of the plane I end up with 35t = 8, or t = 8/35

OpenStudy (anonymous):

ah, so then I just plug it into the parametric form

OpenStudy (amistre64):

3(1+3t) +(1+t) - 5(2-5t) = 2 (3+1-10)+ t(9+1+25) = 2 (-6)+ t(35) = 2 yep; t=8/35

OpenStudy (amistre64):

yep; plug t into the parametric to get the point that is under the other one to measure to :)

OpenStudy (anonymous):

plugging it in I get 59/35, 43/35, -30/35

OpenStudy (amistre64):

im curious; if we take t <normal> and get the magnitude, will it be the same as the distance since we are anchored to 1,1,2 ?

OpenStudy (anonymous):

i'm not sure.

OpenStudy (amistre64):

t tells us how far to stretch the normal to get to any point from our anchored point right?

OpenStudy (anonymous):

just checked the book, and it says the answer is: \[8/\sqrt{35}\]

OpenStudy (anonymous):

for 2a

OpenStudy (anonymous):

so we did something right, and something wrong

OpenStudy (amistre64):

so the normal scaled by t should be the vector from point to point;

OpenStudy (amistre64):

8/35 <3, 1, -5> = sqrt((24)^2+(8)^2+(40)^2)/35 http://www.wolframalpha.com/input/?i=sqrt%28%2824%29%5E2%2B%288%29%5E2%2B%2840%29%5E2%29%2F35 8/sqrt(35)

OpenStudy (amistre64):

t is good, so after finding t is where it might have gone bad ...

OpenStudy (anonymous):

oh alright, so in the future just multiply the t value times the normal?

OpenStudy (amistre64):

IF the point that we are measureing is used to construct out line eqs; then yes

OpenStudy (amistre64):

lets finish the long version tho :) 59/35, 43/35, -30/35 is out point on the plane right?

OpenStudy (anonymous):

i believe so

OpenStudy (amistre64):

(1,1,2) = (35/35, 35/35, 70/35) ( 35, 35, 70)/35 (-59, -43, 30)/35 ---------------- (-16, -8,100)/35 yeah, i think something went bad in the mathing after finding "t"

OpenStudy (amistre64):

so, note to selfs, magnitude t<n> :)

OpenStudy (anonymous):

most likely yes haha

OpenStudy (anonymous):

so i'm really doing t * ||normal||

OpenStudy (anonymous):

rather than just t * normal

OpenStudy (amistre64):

im not sure if t|n| is good; lets see, ive never tried \[\frac{8}{35}|<3,1,-5>|=\frac{8}{35} \sqrt{9+1+25}=\frac{8}{35} \sqrt{35}\] that works too

OpenStudy (anonymous):

how exactly did you do it originally then? when you give the wolfram alpha answer

OpenStudy (amistre64):

scaled n by t then magged it

OpenStudy (anonymous):

so if my professor gives me a similar question on the test, I know how to find t easily, once I find t, what do i do exactly? do I multiply it by the magnitude of the normal?

OpenStudy (amistre64):

\[|\frac{8}{35}<3,1,-5>|=|<\frac{24}{35},\frac{8}{35},-\frac{40}{35}>|\] yes t|n| is good

OpenStudy (amistre64):

lets try the next one :)

OpenStudy (amistre64):

this ones a bit easier since the plane is standard and not tilted

OpenStudy (anonymous):

okay. point = (3, -1, 1) normal is the yz-plane so that's simply <1, 0, 0> if i'm not mistaken

OpenStudy (anonymous):

end up with <3+t, -1, 1>

OpenStudy (anonymous):

t = -3

OpenStudy (anonymous):

-3 ||1^2|| = -3

OpenStudy (amistre64):

yz is the noxplane; how far along the x is our point?

OpenStudy (anonymous):

oh, so yz would be <0, 1, 1>?

OpenStudy (amistre64):

forget the normal for this; there is really no point with the yz plane to find a normal and work out alot of stuff

OpenStudy (amistre64):

think, if a ball is held 5 feet above the floor; how far is it from the floor?

OpenStudy (anonymous):

5 feet

OpenStudy (amistre64):

right; yz is our floor; and x is our height

OpenStudy (amistre64):

whats our height then?

OpenStudy (anonymous):

distance between a point and the yz-plane. so the height is just 3

OpenStudy (amistre64):

exactly :)

OpenStudy (anonymous):

awesome. have time for another, similar, question?

OpenStudy (anonymous):

distance between (-1, 3, 2) and 2x-4y+z=1

OpenStudy (amistre64):

good, normal anchored to our point

OpenStudy (amistre64):

or you can simply enter those parametrics into the plane to get the t

OpenStudy (anonymous):

normal = <2, -4, 1>

OpenStudy (amistre64):

2x-4y+z double them t(4-8+2) t(-2) 2x-4y+z use the point give as -1 3 2 ------- (-2-12+2) = -12 -12+t(-2) = 1 t = -13/2

OpenStudy (anonymous):

<-1 + 2t, 3- 4t,1+t> 2(-1+2t)-4(3-4t)+(1+t) = 1 4t + 16t + t = 14 t = 14/21

OpenStudy (anonymous):

okay so i did something wrong here

OpenStudy (anonymous):

i feel like that is totally different from what I did on 2a, even though they are similar questions

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

i mean, why did I bother with the parametric form for 2a? is that not required?

OpenStudy (amistre64):

-1,3,2 NOT -1,3,1 ^^

OpenStudy (amistre64):

its required, i just shortcuted the process

OpenStudy (anonymous):

ahh

OpenStudy (amistre64):

pull the normal and create parametrics that enter back into the equation so the normal just gets ..... ahh, i messed it up to. square them, not double lol

OpenStudy (anonymous):

the book says the answer is: \[13/\sqrt{21}\]

OpenStudy (anonymous):

i know t = 13 / 21

OpenStudy (amistre64):

2x-4y+z SQUARE the normal t(4+16+1) t(21) 2x-4y+z use the point give as -1 3 2 ------- (-2-12+2) = -12 -12+t(21) = 1 t = 13/21 maybe?

OpenStudy (anonymous):

where is that sq root on the denominator coming from? it was the same for 2a

OpenStudy (amistre64):

its a consequence of form

OpenStudy (amistre64):

|n| = ....21 duh, i got that when i squared and added to begin with sqrt(21) soo \[\frac{3}{21}\sqrt{21}\]

OpenStudy (amistre64):

tell me; is \[\frac{1}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

i don't know. yes?

OpenStudy (anonymous):

yes they are equal ( i used my calculator : ) )

OpenStudy (amistre64):

\[\frac{1}{\sqrt{2}}*1 =\frac{\sqrt{2}}{2}\] \[\frac{1}{\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}} =\frac{\sqrt{2}}{2}\] \[\frac{\sqrt{2}}{\sqrt{2*2}} =\frac{\sqrt{2}}{2}\] \[\frac{\sqrt{2}}{2} =\frac{\sqrt{2}}{2}\]

OpenStudy (amistre64):

its a result of form, not value ....

OpenStudy (anonymous):

ahhhh i see

OpenStudy (amistre64):

some books dont want radicals underneath; others dont care

OpenStudy (anonymous):

\[13/21 \neq 13/\sqrt{21}\]

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

\[\frac{13}{\sqrt{21}}=\frac{3\sqrt{21}}{21}\]

OpenStudy (anonymous):

i understand the form difference, but the fact remains. so where again is that radical coming from? i feel as though i'm missing a step..

OpenStudy (amistre64):

we know t=13/21 right? and |n| = sqrt(sum of squared parts)

OpenStudy (amistre64):

|n| = sqrt(21)

OpenStudy (anonymous):

ohhhh i see

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

ready for another question? Compute the area of the parallelogram spanned by the following vectors. A = <3, -2, 4> B = <5, 1, 1>

OpenStudy (amistre64):

sooo, we can shorten this down perhaps bay saying: the distance can be configured as\[something*\frac{|n|}{|n|^2}\] lol, well at least that will help us eventually

OpenStudy (amistre64):

|axb|

OpenStudy (amistre64):

find the length of the vector produced from A cross B

OpenStudy (anonymous):

in my mind, for distance, it's 1. find t 2. multiply t by the magnitude of the vector

OpenStudy (amistre64):

thats a good thing to have :)

OpenStudy (amistre64):

oh, and once t is found; we can short it to sqrt(denom)

OpenStudy (anonymous):

cool. for the cross product question: i got -6i + 17j + 13k (or <-6, 17, 13>)